Let A & B be 3*3 matrix of real numbers , where A is symmetric & , B is skew symmetric, &(A+B)(A-B)=(A-B)(A+B).if (AB) T =(-1) k AB,where (AB) T is the transpose of the matrix AB, then the possible values of k are Ans=1&3


  • Let A & B be 3*3 matrix of real numbers , where A is symmetric & , B is skew symmetric, &(A+B)(A-B)=(A-B)(A+B).if (AB)T   =(-1)k AB,where  (AB)T  is the transpose of the matrix AB, then the possible values of k are                                                                        

    Ans=1&3



  •  


Grade:12

1 Answers

Pratham Ashish
14 years ago

hi ayush,

(A+B)(A-B)=(A-B)(A+B)

we will get AB = BA

 

use dis property to get one of d values

(AB)T = BTAT

so apply dis n write BT = -B (as dis is a skew symm)

u will get(AB)T   =(-1) AB     (AB= BA , using d given condition)

 

thus getting value of k=1